If change in energy =3×10^-8J h=6.64×10^-34j-s and c = 3×10^8m/s then walength of light is 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
energy change , ∆E = hc/lambda where h is plank's constant , c is velocity of light and  is wavelength.
here h = 6.64 × 10^-34 Js
c = 3 × 10^8 m/s and ∆E = 3 × 10^-18 J 
now, 3 × 10^-18 = 6.64 × 10^-34 × 3 × 10^8/lambda
or, lambda   = 6.64 × 10^-34 × 10^26
= 6.64 × 10^-8 m 
hence, wavelength of the light is 66.4 nm